OpenStudy (anonymous):

Solve for x. x/2 = -5 x = 5/2 x = -5/2 x = -10

OpenStudy (anonymous):

what separates us from the desired form x=.... ?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

do you know what it means when they say "solve for x"

OpenStudy (anonymous):

yes. find what goes there

OpenStudy (anonymous):

so they want an equation that looks like this: x = ....whatever....

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

ok so whats wrong with the equation we are given?

OpenStudy (anonymous):

we don't know what x represents?

OpenStudy (anonymous):

not exactly. we know this though x/2 = -5

OpenStudy (anonymous):

that's some weird form of an x= equation

OpenStudy (anonymous):

so what is wrong with it ?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

you are right that we don't know what x represents. we only know what x/2 (a half of x) represents. and nobody is interested in this.

OpenStudy (anonymous):

we can easily find what x represents though. because if we are given half x, then whole x will just be two times what we have

OpenStudy (anonymous):

so we are bothered by the fraction that has x captured. x/2 = we don't want /2 on the left side. we want only x=

OpenStudy (anonymous):

right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if we do the same operation to the left hand side and to the right hand side the "balance" of the equation is still the same.

OpenStudy (anonymous):

example: x = 3 2x = 6 <this is the same relation or balance

OpenStudy (anonymous):

so because of the balance property of equations, you are always allowed to multiply or divide left hand side and right hand side: it will not change the information the equation carries.

OpenStudy (anonymous):

which one do I do? times or divide?

OpenStudy (anonymous):

x/2 = -5

OpenStudy (anonymous):

which one should we do ?

OpenStudy (anonymous):

ok, do you know how variables work ?

OpenStudy (anonymous):

what is 2 * 2 x

OpenStudy (anonymous):

whattt???

OpenStudy (anonymous):

:) variables

OpenStudy (anonymous):

do you know what is 2*2 =

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

yes. * is multiplication. alright

OpenStudy (anonymous):

so variables are just like a number. the only difference is that they also have the name of the variable.

OpenStudy (anonymous):

and that variable letter never goes away.

OpenStudy (anonymous):

2a + 1 a =

OpenStudy (anonymous):

3a? I don't even know.

OpenStudy (anonymous):

Yes! 3a

OpenStudy (anonymous):

you do what you would always do. and "keep" the letter

OpenStudy (anonymous):

this one is more difficult: 3a + 2 =

OpenStudy (anonymous):

5a?

OpenStudy (anonymous):

yeah, one would think that :) its why I asked in reality, it is just =3a+2 the 2 is not an a. you can not make addition between a variable and a number.

OpenStudy (anonymous):

2a + 9 + 1a =

OpenStudy (anonymous):

Solve for x. = -5 x = 5/2 x = -5/2 x = -10?

OpenStudy (anonymous):

2a + 9 + 1a =

OpenStudy (anonymous):

12

OpenStudy (anonymous):

no: you CAN NOT add variables and numbers

OpenStudy (anonymous):

try again 2a + 9 + 1a =

OpenStudy (anonymous):

2a?

OpenStudy (anonymous):

I don't knowww

OpenStudy (anonymous):

and what about the rest

OpenStudy (anonymous):

2a + 9 + 1a = we can add a terms with each other: 2a+1a=3a we can not add numbers with variables 9 2a + 9 + 1a = 3a +9

OpenStudy (anonymous):

anyway this will get important in a different question. for now, answer me: what is 2*3x

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

hey what is 7* 1/2

OpenStudy (anonymous):

\[7\times\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

you will get really good in algebra

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):



OpenStudy (anonymous):

GOOD

OpenStudy (anonymous):

so in a similar way: \[x \times \frac{ 1 }{ 2 } = \frac{ x }{ 2 }\]

OpenStudy (anonymous):

x/2

OpenStudy (anonymous):

please I need helpSolve for x. = -5 x = 5/2 x = -5/2 x = -10

OpenStudy (anonymous):

how can you cancel this fraction? how can you get rid of it? \[\frac{ 1 }{ 2 } \times x\]

OpenStudy (anonymous):

solve for x \[\frac{ 1 }{ 2 } \times x = -5\] how can we cancel the 1/2 fraction ?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

x/2Solve for x. = -5 x = 5/2 x = -5/2 x = -10

OpenStudy (anonymous):

Okay watch: \[\frac{ 1 }{ 2 } \times 7\]--------------------------\[2 \times \frac{ 1 }{ 2 } \times 7\]--------------------------- \[7\]